Solve the corresponding equation for the appropriate interval by using the following two root finding techniques :

a) algebraic approach



sin(x+ π /2)=ln|x-1|


1
Expert's answer
2022-05-10T18:26:49-0400

"f(x)=sin{(x+\frac{\pi}{2})}"

"f(0)=sin{(0+\frac{\pi}{2})}=sin{\frac{\pi}{2}}=1"

"f(\pi)=sin{(\pi+\frac{\pi}{2})}=sin{\frac{3\pi}{2}}=-1"

"f(\frac{\pi}{2})=sin{(\frac{\pi}{2}+\frac{\pi}{2})}=sin{\pi}=0"

"f(x)=ln|x-1|"

"f(0)=ln|0-1|=ln1=0"

"f(\pi)=ln|\pi-1|=0,76155"

"f(\frac{\pi}{2})=ln|{\frac{\pi}{2}}-1|=-0,56072"

"f'(x)=sin'(x+\frac{\pi}{2})=cos(x+\frac{\pi}{2})"

"f’(x)= cos(x+\frac{\pi}{2})=0"

"x+\frac{\pi}{2}=\frac{\pi}{2}+\pi k, k\in Z"

"x=\frac{\pi}{2}-\frac{\pi}{2}+\pi k, k\in Z"

"x=\pi k,k \in Z"

"f’(x)=(ln|x-1|)’=\frac{1}{|x-1|}"

"\frac{1}{|x-1|}=0"

"x\in R/(1)"